Jamal pulls a 150N sled up a 24° slope at constant speed by a force of 100N. Near the top of hill , he releases the sled. With what acceleration does the sled go down the hill?
1
Expert's answer
2020-09-29T09:41:05-0400

solution

given data

weight of sled (W)=150 N

slope of hill=240

so mass of sled (m)="\frac{150}{9.8}=15.31kg"


when sled will be released then it will come down due to the force of gravity witch is parallel to the slope of hill

so

"150\sin20=ma"

"a" =acceleration of sled when released

"a=\frac{150\times.40}{15.31}=3.9\text m/s^2"

so acceleration of sled will be 3.9m/s^2